OpenStudy (anonymous):

Solve : 6sin^2(x)+2sin^2(2x)=5

OpenStudy (turingtest):

\[6\sin^2x+2\sin^2(2x)=5\]is the problem?

OpenStudy (anonymous):

Yes.

OpenStudy (turingtest):

\[6\sin^2(x)+2\sin^2(2x)=5\]\[3(1-\cos(2x))+(1-\cos(4x))=5\]\[3\cos(2x)+\cos(4x)+1=0\]\[3\cos(2x)+(2\cos^2(2x)-1)+1=0\]\[2\cos^2(2x)+3\cos(2x)=0\]\[\cos (2x)(\cos(2x)+\frac32)=0\]

OpenStudy (anonymous):

Thank you!

OpenStudy (turingtest):

tricky one!

OpenStudy (anonymous):

Indeed .

OpenStudy (anonymous):